Verify cayley hamiltan theorem for A=[ 124212421]

1
Expert's answer
2022-02-01T14:53:41-0500

"A=\begin{pmatrix}
1&2&4\\
2&1&2\\
4&2&1
\end{pmatrix}"

The characteristic polynomial of this matrix is

"P_A(\lambda)=\det(A-\lambda I)=\det\begin{pmatrix}
1-\lambda&2&4\\
2&1-\lambda&2\\
4&2&1-\lambda
\end{pmatrix}="

"(1-\lambda)^3+16+16-16(1-\lambda)-4(1-\lambda)-4(1-\lambda)="

"(1-\lambda)^3-24(1-\lambda)+32"

By the Cayley - Hamilton theorem, it must be "P_A(A)=0". Let's verify this equality.

"P_A(A)=(I-A)^3-24(I-A)+32I"

"I-A=\begin{pmatrix}
0&-2&-4\\
-2&0&-2\\
-4&-2&0
\end{pmatrix}"

"(I-A)^2=\begin{pmatrix}
20&8&4\\
8&8&8\\
4&8&20
\end{pmatrix}"

"(I-A)^3=\begin{pmatrix}
-32 & -48 & -96\\
-48 & -32 & -48\\
-96 & -48 & -32
\end{pmatrix}"

"P_A(A)=\begin{pmatrix}
-32 & -48 & -96\\
-48 & -32 & -48\\
-96 & -48 & -32
\end{pmatrix}-
24\begin{pmatrix}
0&-2&-4\\
-2&0&-2\\
-4&-2&0
\end{pmatrix}+"

"+\begin{pmatrix}
32&0&0\\
0&32&0\\
0&0&32
\end{pmatrix}=\begin{pmatrix}
0&0&0\\
0&0&0\\
0&0&0
\end{pmatrix}"

As we can see, the Cayley - Hamilton equality is true.